What is the solution of y= x+4, and y= -2x+ 1
Answer by stanbon(75887)

What is the solution of
y= x+4,
y= -2x+ 1
-----
x + 4 = -2x+1
3x = 3
x = 1
----
Then y = 1+4 = 5
----
Solution: (1,5)
==================
Cheers,
Stan H.